Two bodies are at temperature 27 degree centigrade and 927 degree centigrade. The heat energy radiated by them will be in the ratio:
A. 1:4
B. 1: 64
C. 1:16
D. 1:256

Grade:12th pass

1 Answers

Arun
4 years ago
According to Stephen's law of radiation energy , 
E  T⁴ 
Here E is the energy of radiation emitted and T is the temperature in Kelvin.
T₁ = 27°C = 27 + 273 = 300 K 
T₂ = 127°C = 927 + 273 = 1200 K 
Now, E₁/E₂ = T₁⁴/T₂⁴ 
E₁/E₂ = (1200/300)² = 256/1 
Hence, the ratio of energies of radiation emitted is 1: 256